# USP: Unified Self-Supervised Pretraining for Image Generation and Understanding

### v2

@@ Abstract
Recent studies have highlighted the interplay between diffusion models and representation learning. Intermediate representations from diffusion models can be leveraged for downstream visual tasks, while self-supervised vision models can enhance the convergence and generation quality of diffusion models. However, transferring pretrained weights from vision models to diffusion models is challenging due to input mismatches and the use of latent spaces. To address these challenges, we propose Unified Self-supervised Pretraining (USP), a framework that initializes diffusion models via masked latent modeling in a Variational Autoencoder (VAE) latent space. USP achieves comparable performance in understanding tasks while significantly improving the convergence speed and generation quality of diffusion models. Our code is available on https://github.com/AMAP-ML/USP.
@@ 2.1 Motivations
Based on the above discussions, we propose a multi-faceted strategy to address the identified challenges. We project the image into its latent space using the frozen off-the-shell VAE used in generative models to address the first half of C2, which can benefit from the advantage of P4. Moreover, we can address the remaining issue of C2 by using curated adaptations in P3. Instead of solving C3 directly, we take an immediate alternative. Drawing inspiration from P2, we design our training objective to foster robust discriminative representations, essential for both image understanding and generation tasks. This approach effectively circumvents the challenges posed by divergent loss functions and labels, thereby harmonizing the training process across both domains. Combining P1, strong representations can greatly alleviate the influence of C1. Furthermore, we carefully calibrate the normalization to increase the matching degree of input data distribution.
@@ 2.2 Model Architecture
Given an input image $I\in\mathbb{R}^{H\times W\times 3}$, its latent feature $I_{vae}$ is obtained through a frozen VAE. For example, using the SD-VAE [71] with $H=W=224$, we obtain a compressed feature $I_{vae}\in\mathbb{R}^{\frac{H}{8}\times\frac{W}{8}\times 4}$. This latent feature is then divided into non-overlapping patches using 2$\times$2 convolutions (referred to as PatchConv), resulting in patches of size $\frac{H}{16}\times\frac{W}{16}\times h$, comparable to those used in [27]. A class token is added to the image tokens and enhanced with their sine-cosine position encodings. These patches are randomly masked with a specified mask ratio $m$. The unmasked patches are first processed by a standard Vision Transformer (ViT) encoder [27] with $h$ hidden dimensions, and then concatenated with the masked patches as input to an asymmetric vision decoder for reconstruction. After pretraining, the decoder is discarded, and only the ViT encoder and PatchConv are retained. Finally, the weights of the pretrained model are used to initialize downstream tasks such as classification, semantic segmentation, and generation. Therefore, this approach doesn’t introduce extra either training cost or memory overhead for downstream tasks.
@@ 3.1 Pretraining
To ensure a fair comparison with MAE [39], we pretrain our models at a resolution of 224$\times$224, although experiments indicate that higher performance can be achieved with larger resolutions. We utilize a mask ratio of 0.75 and apply patch-wise normalization loss as MAE [39]. The VAE is kept frozen throughout this paper. To maintain comparable training efficiency with MAE, we use the VAE to generate image latent features and construct caches in an off-the-shelf manner. Note that we adopt a different image normalization setting, with mean=[0.5, 0.5, 0.5] and std=[0.5, 0.5, 0.5], instead of the default ImageNet normalization.
@@ 3.4 Ablation Study
We perform ablation study in this section and provide additional ablation studies in Section A (appendix).
@@
Additional Adaption for LayerNorm. The results are presented in Table 8(b). Here, “+ LN” denotes that the learnable parameters in LN are utilized, thereby incorporating additional pretrained weights.
@@
Input resolution. We further evaluate our approach at an input resolution of 256$\times$256, consistent with the DiT model’s default configuration. All models are pretrained for 800 epochs. The results, summarized in Table 8(e), demonstrate that higher resolution leads to improved performance across all tasks. However, to ensure fair comparisons with prior work, we maintain the same training budget and use 224×224 as the default input size.
@@
Number of Layers for Weight Initialization. DDAE [88] demonstrated that the most discriminative features in diffusion models reside in the intermediate layers. Similarly, REPA [97] showed that applying representation alignment loss to the first eight layers improves performance. These findings indicate that early layers primarily capture understanding, while later layers focus on generation. Consequently, as detailed in Table 8(g), we reinitialize the pretrained weights of the last 2, 4, and 6 layers with random weights. Although increasing the number of randomly initialized layers deteriorates generation quality, the performance still exceeds that of the baseline, highlighting the critical importance of pretraining each layer in USP.
@@ 3.5 Discussion
VAE for image understanding. From the efficient architecture perspective, it’s not a perfect idea to apply VAE on designing image backbones for classification. After all, the philosophy of VAE is to pursue higher compression while maintaining stronger reconstruction. The raw image itself is lossless and more efficient. However, it’s interesting to see that given a latent representation encoded by a good VAE, we can achieve at least comparable recognition performance.
@@ 4 Related work
Self supervised learning. Self-supervised learning (SSL) leverages unlabeled data to learn visual representations via pretext tasks. Contrastive learning methods, such as SimCLR [13] and MoCo [38], contrast positive pairs with negative samples from mini-batches, enhanced by data augmentations. Self-distillation methods like BYOL [33] and SimSiam [14] achieve strong performance without negative samples, aligning representations through self-distillation. Masked Image Modeling (MIM) methods [39, 4, 67, 84, 83, 11] extend denoising autoencoders [82], recovering masked regions in images.
@@
Accelerating Convergence of Diffusion Transformers. Our focus is on improving diffusion models without largely altering the architecture, unlike approaches such as [19, 95], which introduce architectural changes to enhance performance. Given this limitation, the core idea of accelerating the convergence of diffusion models is to leverage visual representations [30, 101, 97, 95, 93]. Most approaches achieve this by adding various types of losses during the training of diffusion models [101, 97]. For instance, REPA [97] aligns the representations of diffusion models with external visual representations, enhancing both convergence speed and generation quality. Similarly, studies such as VA-VAE [95] and ReaLs et al. [93] propose aligning representations during VAE training to address the lack of clear semantic representations in traditional VAE latents. MaskDiT [101] adds an MAE decoder to DiT and incorporates a masked token reconstruction objective during training, improving the training efficiency of diffusion models. MDT [30] performs reconstruction on all image tokens and introduces carefully designed modules to the original DiT architecture, achieving better generation performance despite the increased training cost. While these works focus on accelerating the convergence and improving the performance of diffusion models, our method decouples pretraining from downstream tasks such as image generation and recognition. This allows our approach to not only achieve significant improvements in generation tasks but also enhance performance on understanding tasks, offering a more unified and efficient solution.
@@ 5 Conclusion
In this work, we delve into the pretraining-fine-tuning paradigm, targeting both image recognition and generation tasks. By conducting masked latent modeling within the latent space of a Variational Autoencoder (VAE), we achieve the learning of unified, robust representations. This approach necessitates only a single training phase, after which the model can be seamlessly integrated into downstream tasks via weight initialization, thereby yielding substantial performance improvements. Notably, our method incurs no additional training costs or GPU memory overhead for downstream tasks. We posit that our approach serves as a potent baseline for the research community.
@@ Appendix A More Ablation Studies
High-Resolution Results Table 10 demonstrates the effectiveness of USP at a higher resolution of 512×512. We initialize the downstream image generation task using weights obtained from pretraining with USP at 256 resolution and directly transfer them to the 512 resolution setting, adjusting the positional encodings via bilinear interpolation. The results confirm that USP maintains strong performance and transferability under higher-resolution settings.
@@
Table 10: Results at 512×512 Resolution.
Image Normalization (IN). Image normalization is a standard transformation in the community, and the default mean ([0.485, 0.456, 0.406]) and std ([0.229, 0.224, 0.225]) are widely used. However, VAE of DiT [66] has a different setting (mean=[0.5, 0.5, 0.5], std=[0.5, 0.5, 0.5]). We perform pretraining using these two groups and report the downstream results in Table 11. Although the default setting of ImageNet has lower loss (0.375), it doesn’t bring in higher performance. Therefore, we utilize the settings of SD-VAE [71].
@@ Appendix B More Related Work
MLLMs for Unified Understanding and Generation Multimodal Large Language Models (MLLMs) have recently drawn extensive attention from both academia and industry. MLLMs [2, 16, 60] enable visual question answering in multimodal understanding tasks by aligning image embeddings with textual embeddings and jointly feeding them into a large language model, ultimately yielding text token IDs.
@@ C.2 Fully Tuning on ImageNet
The hyperparameter setting for fine-tuning in ImageNet is shown in 14.
@@ C.3 Image Generation on ImageNet
The hyperparameter setting for generation in ImageNet is shown in 14.
@@ C.4 Image Generation
We visualize more image generation results from DiT-XL/2 and SiT-XL/2, as shown in Figure 7 and Figure 8, respectively. All results are generated with a classifier-free guidance (CFG) scale of 4.0.

### v3

@@ Abstract
Recent studies have highlighted the interplay between diffusion models and representation learning. Intermediate representations from diffusion models can be leveraged for downstream visual tasks, while self-supervised vision models can enhance the convergence and generation quality of diffusion models. However, transferring pretrained weights from vision models to diffusion models is challenging due to input mismatches and the use of latent spaces. To address these challenges, we propose Unified Self-supervised Pretraining (USP), a framework that initializes diffusion models via masked latent modeling in a Variational Autoencoder (VAE) latent space. USP achieves comparable performance in understanding tasks while significantly accelerating convergence and enhancing generation quality in diffusion models. Our code is available on https://github.com/AMAP-ML/USP.
@@ 2.1 Motivations
Based on the above discussions, we propose a multi-faceted strategy to address the identified challenges. We project the image into its latent space using the frozen off-the-shelf VAE used in generative models to address the first half of C2, which can benefit from the advantage of P4. Moreover, we can address the remaining issue of C2 by using curated adaptations in P3. Instead of solving C3 directly, we take an immediate alternative. Drawing inspiration from P2, we design our training objective to foster robust discriminative representations, essential for both image understanding and generation tasks. This approach effectively circumvents the challenges posed by divergent loss functions and labels, thereby harmonizing the training process across both domains. Combining P1, strong representations can greatly alleviate the influence of C1. Furthermore, we carefully calibrate the normalization to increase the matching degree of input data distribution.
@@ 2.2 Model Architecture
Given an input image $I\in\mathbb{R}^{H\times W\times 3}$, its latent feature $I_{vae}$ is obtained through a frozen VAE. For example, using the SD-VAE [71] with $H=W=224$, we obtain a compressed feature $I_{vae}\in\mathbb{R}^{\frac{H}{8}\times\frac{W}{8}\times 4}$. This latent feature is then divided into non-overlapping patches using 2$\times$2 convolutions (referred to as PatchConv), resulting in patches of size $\frac{H}{16}\times\frac{W}{16}\times h$, comparable to those used in [27]. A class token is added to the image tokens and enhanced with their sine-cosine position encodings. These patches are randomly masked with a specified mask ratio $m$. The unmasked patches are first processed by a standard ViT encoder [27] with $h$ hidden dimensions, and then concatenated with the masked patches as input to an asymmetric vision decoder for reconstruction. After pretraining, the decoder is discarded, and only the ViT encoder and PatchConv are retained. Finally, the weights of the pretrained model are used to initialize downstream tasks such as classification, semantic segmentation, and generation. Therefore, this approach doesn’t introduce either extra training cost or memory overhead for downstream tasks.
@@ 3.1 Pretraining
To ensure a fair comparison with MAE [39], we pretrain at a resolution of 224$\times$224, although experiments indicate that higher performance can be achieved with larger resolutions. We utilize a mask ratio of 0.75 and apply patch-wise normalization loss as MAE [39]. The VAE is kept frozen throughout this paper. To maintain comparable training efficiency with MAE, we use the VAE to generate image latent features and construct caches in an off-the-shelf manner. Note that we adopt a different image normalization setting, with mean=[0.5, 0.5, 0.5] and std=[0.5, 0.5, 0.5], instead of the default ImageNet normalization.
@@ 3.4 Ablation Study
In this section, we conduct an ablation study, with further details provided in Section A (appendix).
@@
Additional Adaptation for LayerNorm. The results are presented in Table 8(b). Here, “+ LN” denotes that the learnable parameters in LN are utilized, thereby incorporating additional pretrained weights.
@@
Input Resolution. We further evaluate our approach at an input resolution of 256$\times$256, consistent with the DiT model’s default configuration. All models are pretrained for 800 epochs. The results, summarized in Table 8(e), demonstrate that higher resolution leads to improved performance across all tasks. However, to ensure fair comparisons with prior work, we maintain the same training budget and use 224$\times$224 as the default input size.
@@
Number of Layers for Weight Initialization. DDAE [88] demonstrated that the most discriminative features in diffusion models reside in the intermediate layers. Similarly, REPA [97] showed that applying representation alignment loss to the first 8 layers improves performance. These findings indicate that early layers primarily capture understanding, while later layers focus on generation. Consequently, as detailed in Table 8(g), we reinitialize the pretrained weights of the last 2, 4, and 6 layers with random weights. Although increasing the number of randomly initialized layers deteriorates generation quality, the performance still exceeds that of the baseline, highlighting the critical importance of pretraining each layer in USP.
@@ 3.5 Discussion
VAE for Image Understanding. From the efficient architecture perspective, it’s not a perfect idea to apply VAE to designing image backbones for classification. After all, the philosophy of VAE is to pursue higher compression while maintaining stronger reconstruction. The raw image itself is lossless and more efficient. However, it’s interesting to see that given a latent representation encoded by a good VAE, we can achieve at least comparable recognition performance.
@@ 4 Related work
Self-Supervised Learning. Self-supervised learning (SSL) leverages unlabeled data to learn visual representations via pretext tasks. SimCLR [13] and MoCo [38], contrast positive pairs with negative samples from mini-batches, enhanced by data augmentations. Self-distillation methods like BYOL [33] and SimSiam [14] achieve strong performance without negative samples, aligning representations through self-distillation. Masked Image Modeling (MIM) methods [39, 4, 67, 84, 83, 11] extend denoising autoencoders [82], recovering masked regions in images.
@@
Accelerating Convergence of Diffusion Transformers. Our focus is on improving diffusion models without largely altering the architecture, unlike approaches such as [19, 95], which introduce architectural changes to enhance performance. Given this limitation, the core idea of accelerating the convergence of diffusion models is to leverage visual representations [30, 101, 97, 95, 93]. Most approaches achieve this by adding various types of losses during the training of diffusion models [101, 97]. For instance, REPA [97] aligns the representations of diffusion models with external visual representations, enhancing both convergence speed and generation quality. Similarly, studies such as VA-VAE [95] and ReaLs et al. [93] propose aligning representations during VAE training to address the lack of clear semantic representations in traditional VAE latents. MaskDiT [101] adds an MAE decoder to DiT and incorporates a masked token reconstruction objective during training, improving the training efficiency of diffusion models. MDT [30] performs reconstruction on all image tokens and introduces carefully designed modules to the original DiT architecture, achieving better generation performance despite the increased training cost. While these works focus on accelerating the convergence and improving the performance of diffusion models, our method decouples pretraining from downstream tasks such as image generation and recognition. This allows our approach to offer a more unified and efficient solution.
@@ 5 Conclusion
In this work, we delve into the pretraining-fine-tuning paradigm, targeting both image recognition and generation tasks. By conducting masked latent modeling within the latent space of a VAE, we achieve the learning of unified, robust representations. This approach necessitates only a single training phase, after which the model can be seamlessly integrated into downstream tasks via weight initialization, thereby yielding substantial performance improvements. Notably, our method incurs no additional training costs or GPU memory overhead for downstream tasks. We posit that our approach serves as a potent baseline for the research community.
@@ Appendix A More Ablation Studies
High-Resolution Results Table 10 demonstrates the effectiveness of USP at a higher resolution of 512$\times$512. We initialize the downstream image generation task using weights obtained from pretraining with USP at 256 resolution and directly transfer them to the 512 resolution setting, adjusting the positional encodings via bilinear interpolation. The results confirm that USP maintains strong performance and transferability under higher-resolution settings.
@@
Table 10: Results at 512$\times$512 Resolution.
Image Normalization (IN). Image normalization is a standard transformation in the community, and the default mean ([0.485, 0.456, 0.406]) and std ([0.229, 0.224, 0.225]) are widely used. However, VAE of DiT [66] has a different setting (mean=[0.5, 0.5, 0.5], std=[0.5, 0.5, 0.5]). We perform pretraining using these two groups and report the downstream results in Table 11. Although the default setting of ImageNet has a lower loss (0.375), it doesn’t bring in higher performance. Therefore, we utilize the settings of SD-VAE [71].
@@ Appendix B More Related Work
MLLMs for Unified Understanding and Generation. Multimodal Large Language Models (MLLMs) have recently drawn extensive attention from both academia and industry. MLLMs [2, 16, 60] enable visual question answering in multimodal understanding tasks by aligning image embeddings with textual embeddings and jointly feeding them into a large language model, ultimately yielding text token IDs.
@@ C.2 Fully Tuning on ImageNet
The hyperparameter setting for fine-tuning in ImageNet is shown in Table 13.
@@ C.3 Image Generation on ImageNet
The hyperparameter setting for generation in ImageNet is shown in Table 14.
@@ C.4 Image Generation
We visualize more image generation results from DiT-XL/2 and SiT-XL/2, as shown in Figure 7 and Figure 8, respectively. All results are generated with a CFG scale of 4.0.